OpenStudy (anonymous):

Determine which function is a solution to the differential equation xy ' + 2y = 0.

OpenStudy (anonymous):

xy'+ 2y = 0 x = 2y/y' \[\int\limits x = \int\limits \frac{ 2y }{ y' }\] not sure what to do after this part

OpenStudy (zzr0ck3r):

This looks like pug and play, did they give you some functions to try?

OpenStudy (zzr0ck3r):

plug*

OpenStudy (usukidoll):

I think this equation is separable.